OpenStudy (anonymous):

Jack has type A blood. His father was type AB and his mother was type O. Jill is type B. Her father was type B, and her mother was type AB. a. What possible offspring (in relation to blood type) could Jack and Jill have? (1 mark) a. A, B, and 0 b. A and B c. A, B, and AB d)A,B,0, AND AB e)0,A,and AB show on a punnet square.

OpenStudy (anonymous):

i would prolly say B. A and B blood types

OpenStudy (anonymous):

how did u get that ?i would have to make 2 punnet squares hey?for jack and jills parents aye?

OpenStudy (anonymous):

or do i find jack and jill hey?

OpenStudy (anonymous):

@thomaster do u understand how he got this???

OpenStudy (anonymous):

its obvious when you look at the parent blood types and just do prosses of elimenation

thomaster (thomaster):

Yes. In blood types you have 3 types of alleles: \(\large\sf I^A\), \(\large\sf I^B\) and \(\large\sf i\) In a genotype you have a set of 2 alleles which convert to blood types as follows: \(\large\sf I^AI^A\) = blood type A \(\large\sf I^Ai\) = blood type A \(\large\sf I^BI^B\) = blood type B \(\large\sf I^Bi\) = blood type B \(\large\sf I^AI^B\) = blood type AB \(\large\sf ii\) = blood type 0 Jack's father = \(\large\sf I^AI^B\) Jack's mother = \(\large\sf ii\) Jack will get 1 allele from each parent. Since he has blood type A, he got \(\large\sf I^A\) from his father and \(\large\sf i\) from his mother which means Jack has genotype \(\large\sf I^Ai\) Jill's father = \(\large\sf I^BI^B\) or \(\large\sf I^Bi\) Jill's mother = \(\large\sf I^AI^B\) So Jill's genotype is \(\large\sf I^BI^B\) or \(\large\sf I^Bi\) It's sure she got the \(\large\sf I^B\) from her mother but not sure if she got the \(\large\sf I^B\) from her father or the \(\large\sf i\) Now you combine Jack and Jill's genotypes in 2 different Punnett squares since Jill has 2 possible genotypes.

thomaster (thomaster):